# Efficient Reinforcemen Learning via Decoupling Exploration and Utilization

### v3

@@ V-C Evaluation of OPARL in DMControl Environment
Table I delineates the mean performance metrics of the algorithm after a training regimen spanning one million time steps. The empirical outcomes significantly indicate that the OPARL algorithm surpasses the comparative algorithms in the majority of the environments tested (14 out of 18). Specifically, when juxtaposed with the REDQ, SAC, and TD3 algorithms, OPARL demonstrated performance enhancements of 14.30%, 4.08%, and 7.36%, respectively. Notably, in the ”swimmer-swimmer6” and ”swimmer-swimmer15” scenarios, OPARL’s performance increments were particularly pronounced, amounting to 71.04% and 100.59%, respectively. A more granular analysis of the learning curves, commencing from a foundational level, is provided in the appendix A. These findings not only underscore the adaptability and superiority of OPARL across a diverse array of environments but also lay a solid groundwork for further exploration of its potential in complex task applications.
V-D Ablation Study
@@
A series of ablation studies were conducted to assess the impact of novel components integrated into our algorithm’s structure. The experimental design involved the independent application of an optimistic exploration strategy—selecting the maximal value from an ensemble of five values—and a pessimistic exploration strategy—selecting the minimal value from the ensemble. Additionally, we examined the effect of not introducing extra ensemble values, thus maintaining the original configuration with only two values. The evaluation results are presented in the figure. While Yarats et al.’s study[54] suggests that adequate exploration and state coverage can enable standard reinforcement learning to achieve equivalent outcomes without a pessimistic strategy, both prior work on decoupled strategy learning[55] and our experimental findings consistently demonstrate that to enhance performance, optimistic exploration must be coupled with pessimistic constraints.
VI CONCLUSION
@@
Appendix A Performance Comparison Diagram of Each Model in DMControl Environment.
Our algorithm demonstrates superior performance compared to the original TD3 model across all test environments, with particularly notable achievements in the Swimmer swimmer6 and Swimmer swimmer15 scenarios. In these specific environments, our model exhibits significantly enhanced performance when contrasted with the original TD3 model.
Appendix B Detailed Description of The Environment
Ant: The environment is modeled after the one introduced by Schulman, Moritz, Levine, Jordan, and Abbeel in their work[56]. The ant model is a quadruped robot, featuring a freely rotating torso with four attached legs, each leg comprising two links. The primary objective is to achieve coordinated forward (rightward) movement by applying torques to eight hinges that connect the leg links and the torso, encompassing nine parts and eight hinges in total. This environment allows versatile movement simulations within its framework. The observation space includes positional and velocity information for various parts of the ant’s body. Rewards in this environment are structured around maintaining a healthy state, moving forward, and minimizing control costs.
Hopper:The environment is predicated on the foundational work of Erez, Tassa, and Todorov in ”Infinite Horizon Model Predictive Control for Nonlinear Periodic Tasks.” The aim of this environment is to augment the quantity of independent state and control variables in contrast to classical control settings. The hopper robot is depicted as a two-dimensional monopodal entity, comprising four primary bodily segments: an upper torso, a thigh situated midsection, a lower leg beneath the thigh, and a solitary foot that serves as the foundational support for the entire structure. The objective is to execute forward (rightward) hopping maneuvers by exerting torques upon the three hinges that interconnect the quadruple body segments. The model’s overarching goal is to facilitate effective hopping and balance, whilst simultaneously optimizing the distance progressed forward and forestalling any potential collapse.
Walker2d: The environment extends the hopping robot scenario introduced by Erez, Tassa, and Todorov in ”Infinite Horizon Model Predictive Control for Nonlinear Periodic Tasks” by incorporating an additional set of limbs, thereby enabling the robot to achieve forward locomotion rather than mere hopping. This adaptation aligns with the broader aim of MuJoCo-based environments to expand the repertoire of independent state and control variables relative to classical control environments. The Walker robot is conceptualized as a bi-dimensional bipedal entity, structurally segmented into four principal components: a singular torso serving as the upper segment from which two legs diverge; two thighs situated below the torso; two lower legs beneath the thighs; and a pair of feet affixed to the lower legs, forming the base for the entire structure. The primary objective is the calibrated manipulation of both pairs of feet, legs, and thighs, employing torques across six hinges interlinking the sextet of body segments, to facilitate a concerted forward (rightward) progression. The Walker model shares similarities with the Hopper in terms of its action space, which comprises torques applied across joints, with the overarching goal of accomplishing steady bipedal traversal and maximizing the distance covered in the forward trajectory.
Humanoid:The environment is predicated on the groundwork laid by Tassa, Erez, and Todorov in ”Synthesis and Stabilization of Complex Behaviors through Online Trajectory Optimization.” The Humanoid model is an intricate 3D bipedal robot, devised to emulate human locomotion and kinetics. It comprises a torso (abdomen) with appendages articulated as a pair of legs and arms. Each limb bifurcates into two segments, akin to human knees and elbows. The principal aim of the environment is to navigate forward at the highest possible velocity while maintaining equilibrium. The action space encompasses the modulation of torques across multiple bodily articulations, while the observation space encompasses the positional and velocity data of the respective body segments. The model presents the challenge of executing efficient bipedal ambulation and a repertoire of other complex maneuvers.
Appendix C Detailed Description of The Comparison Algorithm
Appendix D The Specific Parameter Setting of OPARL Algorithm
| Hyperparameter | Value |
| Qnum(The number of ensemble Q’s) | 5 |
| start-timesteps(Time steps initial random policy is used) | 25000 |
| max-timesteps(Max time steps to run environment) | 1000000 |
| expl-noise(Std of Gaussian exploration noise) | 0.1 |
| discount(Discount factor) | 0.99 |
| tau(Target network update rate) | 0.005 |
| policy-noise(Noise added to target policy during critic update) | 0.2 |
| noise-clip(Range to clip target policy noise) | 0.5 |
| policy-freq(Frequency of delayed policy updates) | 2 |
| policy-utd(The ratio of the number of training sessions to the number of interactions with the environment) | 1 |
| policy-interval(The number of interactions with the environment per exploration) | 1 |
| N(Enter the number of actions generated by a state) | 1 |
| bili(The ratio of optimistic exploration to pessimistic exploration) | 2 |

### v4

@@ 5.3 Evaluation of OPARL in DMControl Environment
Table 1 delineates the mean performance metrics of the algorithm after a training regimen spanning one million time steps. The empirical outcomes significantly indicate that the OPARL algorithm surpasses the comparative algorithms in the majority of the environments tested (14 out of 18). Specifically, when juxtaposed with the REDQ, SAC, and TD3 algorithms, OPARL demonstrated performance enhancements of 14.30%, 4.08%, and 7.36%, respectively. Notably, in the ”swimmer-swimmer6” and ”swimmer-swimmer15” scenarios, OPARL’s performance increments were particularly pronounced, amounting to 71.04% and 100.59%, respectively. These findings not only underscore the adaptability and superiority of OPARL across a diverse array of environments but also lay a solid groundwork for further exploration of its potential in complex task applications.
5.4 Ablation Study
@@
Figure 3 presents a series of ablation studies aimed at evaluating the impact of novel components integrated into the architecture of our algorithm. The experimental design involved the independent application of an optimistic exploration strategy—selecting the maximal value from an ensemble of five values—and a pessimistic exploration strategy—selecting the minimal value from the ensemble. Additionally, we examined the effect of not introducing extra ensemble values, thus maintaining the original configuration with only two values. The evaluation results are presented in the figure. While Yarats et al.’s study[55] suggests that adequate exploration and state coverage can enable standard reinforcement learning to achieve equivalent outcomes without a pessimistic strategy, both prior work on decoupled strategy learning[56] and our experimental findings consistently demonstrate that to enhance performance, optimistic exploration must be coupled with pessimistic constraints.
6 CONCLUSION